OpenStudy (anonymous):

Each day, Karen makes 45 loaves of bread at the bakery where she works. Each loaf of bread requires 5 cups of flour. One pound of flour is equivalent to about 4 cups of flour. About how many pounds of flour does Karen need in order to make the loaves of bread? A. 900 B. 45 C. 225 D. 56.25

ganeshie8 (ganeshie8):

\(\large \frac{45 ~loaves }{1~day} \times \frac{5 ~cups~flour}{1~loave} \times \frac{1 ~pound}{4~cups ~flour}\)